### v3

# L-equivalence and Fourier–Mukai partners of cubic fourfolds

Abstract.
We study $\mathbb{L}$-equivalence in the Grothendieck ring of varieties and its interaction with categorical invariants of cubic fourfolds. Assuming a Derived Torelli-type criterion for Kuznetsov components and a mild condition on the discriminant of the transcendental lattice, we prove a counting formula for Fourier–Mukai partners of such cubic fourfolds. As an application, we exhibit cubic fourfolds with a fixed algebraic lattice admitting a unique non-trivial Fourier–Mukai partner, which is trivially $\mathbb{L}$-equivalent to the original. Finally, we show that $\mathbb{L}$-equivalence classes of cubic fourfolds are finite.
1. Introduction
The Grothendieck ring of varieties over a field $k$ is the quotient of the free abelian group generated by isomorphism classes of varieties over $k$, by the subgroup generated by expressions of the form
$X]-[Z]=[X\setminus Z],$
where $Z\subset X$ is a closed subvariety. Multiplication in $K_{0}(\operatorname{Var}_{k})$ is defined by the fibre product over $k$:
$X]\cdot[Y]=[X\times_{k}Y].$
The class $\mathbb{L}\coloneqq[\mathbb{A}^{1}]\in K_{0}(\operatorname{Var}_{k})$ is called the Lefschetz motive. The Lefschetz motive is a zero-divisor in $K_{0}(\operatorname{Var}_{\mathbb{C}})$, as was proved by Borisov [6]. Following this observation, Kuznetsov and Shinder defined $\mathbb{L}$-equivalence:
Definition 1.1 ([31]).
Two varieties $X,Y$ over a field $k$ are $\mathbb{L}$-equivalent, denoted $X\sim_{\mathbb{L}}Y$, if there exists a $n\geq 0$ such that
$\mathbb{L}^{n}([X]-[Y])=0\in K_{0}(\operatorname{Var}_{k}).$
$\mathbb{L}$-equivalence seems to be closely related to derived equivalence, especially for projective hyperkähler manifolds. At the time of writing, many examples of $\mathbb{L}$-equivalent varieties have been constructed. In the case of K3 surfaces, see [31, 18, 26, 28] and [53], also related with $\mathbb{L}$-equivalent degree five elliptic curves. In higher dimensions, $\mathbb{L}$-equivalent Calabi-Yau varieties are considered in [6, 37, 50, 27, 29, 36, 7]. Examples of $\mathbb{L}$-equivalence for hyperkähler manifolds appear in [47, 4], for Gushel-Mukai surfaces in [52], for cubic fourfolds in [13].
All currently known examples of $\mathbb{L}$-equivalent hyperkähler manifolds are also derived equivalent. In fact, for very general K3 surfaces, $\mathbb{L}$-equivalence implies derived equivalence [40]. On the other hand, there exist examples of derived equivalent hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type that are not $\mathbb{L}$-equivalent, for any $n\geq 2$ [41]. Notably, there are no known examples of K3 surfaces that are derived equivalent but not $\mathbb{L}$-equivalent.
Conjecture 1.2.
[41] If $X$ and $Y$ are $\mathbb{L}$-equivalent projective hyperkähler manifolds, then they are derived equivalent.
In this paper, we study $\mathbb{L}$-equivalence and Fourier–Mukai partnership for cubic fourfolds over the complex numbers. For a cubic fourfold $X$, there exists a canonical semiorthogonal decomposition
$\mathcal{D}^{b}(X)\simeq\langle\mathcal{A}_{X},\mathcal{O}_{X},\mathcal{O}_{X}(1),\mathcal{O}_{X}(2)\rangle.$
The component $\mathcal{A}_{X}$ is called the Kuznetsov component of $\mathcal{D}^{b}(X)$. Since cubic fourfolds are Fano varieties (in the sense that their anti-canonical bundles are ample), it follows from Bondal and Orlov’s Reconstruction Theorem [5] that any Fourier–Mukai equivalence $\mathcal{D}^{b}(X)\simeq\mathcal{D}^{b}(Y)$ between the derived categories of cubic fourfolds induces an isomorphism $X\simeq Y$. However, there are many examples of non-isomorphic cubic fourfolds $X$, $Y$ whose derived categories have equivalent Kuznetsov components $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$. Such cubic fourfolds are called Fourier–Mukai partners.
We therefore propose the following conjecture related to Conjecture 1.2. This conjecture has also recently appeared in [4].
Conjecture 1.3.
If $X$ and $Y$ are $\mathbb{L}$-equivalent cubic fourfolds, then they are Fourier–Mukai partners:
$X\sim_{\mathbb{L}}Y\implies\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$
Remark 1.4.
Conjecture 1.3 should also follow from Kontsevich’s expectation that semiorthogonal decompositions are canonical (up to mutation) and functorial with respect to geometric operations. See, for further references [12, 15].
We note that the $\mathbb{L}$-equivalence relation in the context of cubic fourfolds was first studied by Galkin and Shinder [14]. In that paper, the so-called rational defect $\mathcal{M}_{X}$ of a cubic fourfold $X$ plays a crucial role. The rational defect is defined to be
$\mathcal{M}_{X}\coloneqq\frac{[X]-[\mathbb{P}^{4}]}{\mathbb{L}}\in K_{0}(\operatorname{Var}_{\mathbb{C}})[\mathbb{L}^{-1}].$
If $X$ is rational, the equation $X]=[\mathbb{P}^{4}]+\mathcal{M}_{X}\cdot\mathbb{L}$ holds in the (non-localised) Grothendieck ring $K_{0}(\operatorname{Var}_{\mathbb{C}})$, and $\mathcal{M}_{X}$ is a sum of smooth projective surfaces [14, Lemma 2.1]. In fact, in all of the cases where we can compute $\mathcal{M}_{X}$ for rational cubic fourfolds, we have the equation [31, 26]
(1.1) $X]=1+\mathbb{L}^{2}+\mathbb{L}^{4}+[S]\mathbb{L},$
where $S$ is a K3 surface.
In this paper, we construct pairs of $\mathbb{L}$-equivalent cubic fourfolds $X$, $Y$ whose classes in the Grothendieck ring are of the form (1.1):
Proposition 1.5 (See Proposition 4.5 and Proposition 4.7).
Suppose $X$ is a cubic fourfold with one of the following two properties:
  1. A)
    The primitive part $A(X)_{\mathrm{prim}}$ of the algebraic lattice is isometric to
    $L_{2,9}\coloneqq\left(\begin{matrix}14&2\\ 2&18\end{matrix}\right).$
  2. B)
    The algebraic lattice $A(X)\subset H^{4}(X,\mathbb{Z})$ is isometric to the lattice
    $L_{10}\coloneqq\left(\begin{matrix}3&1&1\\ 1&3&0\\ 1&0&10\end{matrix}\right).$
Moreover, suppose that $X$ is very general with one of these properties. Then $X$ has a non-trivial Fourier–Mukai partner $Y$ for which there exists a unique K3 surface $S$ with the property that
$X]=[Y]=1+\mathbb{L}^{2}+\mathbb{L}^{4}+[S]\mathbb{L}\in K_{0}(\operatorname{Var}_{\mathbb{C}}).$
Finally, if $X$ satisfies Property B, then $Y$ is the unique non-trivial Fourier–Mukai partner of $X$.
The $\mathbb{L}$-equivalent cubic fourfolds in Proposition 1.5 are trivially $\mathbb{L}$-equivalent, in the sense that the motives $[X]$ and $[Y]$ are equal in $K_{0}(\operatorname{Var}_{\mathbb{C}})$. Other examples of trivially $\mathbb{L}$-equivalent cubic fourfolds can be found in [13].
Question 1.6.
Do there exist non-trivially $\mathbb{L}$-equivalent cubic fourfolds?
We prove Proposition 1.5 by constructing a Counting Formula for Fourier–Mukai partners of certain cubic fourfolds. Specifically, we will work with cubic fourfolds satisfying the following Definition.
Definition 1.7.
A cubic fourfold $X$ is said to satisfy Derived Torelli if it has the property that for any cubic fourfold $Y$, the following two statements are equivalent:
  1. (1)
    $X$ and $Y$ are Fourier–Mukai partners, that is, we have an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$.
  2. (2)
    There is a Hodge isometry between transcendental lattices $T(X)\simeq T(Y)$.
Note that $(1)\implies(2)$ always holds, see Remark 3.1. There is a large class of cubic fourfolds which satisfy Derived Torelli, for instance if $X$ is a general cubic fourfold not contained in a Hassett divisor (see also Proposition 3.2 and Remark 3.3).
Inspired by the works [19, 35, 9], we prove the following:
Theorem 1.8 (Theorem 3.5).
Let $X$ be a cubic fourfold which satisfies Derived Torelli. Assume moreover that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
[…]
Here, $N=A(X)_{\text{prim}}$ and the action of $O(N^{\prime})$ on $O(A_{T(X)})$ is the one defined by Remark 3.4, and $\mathcal{H}(N)$ is the set of lattices in the genus of $N$ which contain no vectors of square $2$, nor any vectors of square $6$ and divisibility $3$.
We also prove a partial counting formula for Fourier–Mukai partners of a cubic fourfold $X$ of which the primitive algebraic lattice $N\coloneqq A(X)_{\mathrm{prim}}$ has discriminant indivisible by $3$. For such a cubic fourfold, we prove in Proposition 3.12 that the number of Fourier–Mukai partners $Y$ which satisfy $A(Y)_{\mathrm{prim}}\simeq N^{\prime}\in\mathcal{H}(N)$ is given by
$|O(N^{\prime})\backslash O(A_{N})/O_{\mathrm{Hodge}}(T(X))|.$
Note that $X$ may have Fourier–Mukai partners $Y$ for which the discriminant of $A(Y)_{\mathrm{prim}}$ is a multiple of $3$. In this case we have $A(Y)_{\mathrm{prim}}\notin\mathcal{H}(N)$. We give examples of this phenomenon in Remark 4.8.
Finally, we study the following equivalence relations for cubic fourfolds (we copy the notation from [8]). For two cubic fourfolds $X$, $Y$, we write:
  1. BE:
    $X$ and $Y$ are birational: $X\overset{\simeq}{\dashrightarrow}Y$
  2. BF:
    $X$ and $Y$ have birational Fano varieties of lines: $F(X)\overset{\simeq}{\dashrightarrow}F(Y)$
  3. FM:
    The Kuznetsov components of $X$ and $Y$ are equivalent: $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$
  4. DF:
    The Fano varieties of lines of $X$ and $Y$ are derived equivalent: $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$
  5. LE:
    $X$ and $Y$ are $\mathbb{L}$-equivalent: $X\sim_{\mathbb{L}}Y$
  6. LF:
    The Fano varieties of lines of $X$ and $Y$ are $\mathbb{L}$-equivalent: $F(X)\sim_{\mathbb{L}}F(Y)$
Theorem 1.9 (Theorem 5.3, Theorem 5.4, Proposition 5.6, Corollary 5.9, Theorem 5.7).
For two cubic fourfolds $X$, $Y$, we have the following implications, where the implications marked by $\dagger$ are proved only when $X$ and $Y$ satisfies Derived Torelli and $X$ is such that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$:
[…]
The fact that $\mathbb{L}$-equivalent cubic fourfolds are Fourier–Mukai partners was recently independently proved in [4] for very general cubic fourfolds in $\mathcal{C}_{d}$ for $d\equiv 0,2\pmod{6}$ not divisible by $9$. We note that such cubic fourfolds satisfy Derived Torelli. They also prove LE $\implies$ LF and the fact that BE does not imply LE.
For the convenience of the reader, let us complete Diagram 1.2 with the implications that are already known. The implication FM $\implies$ DF was proved recently in [30], and the fact that BF $\implies$ DF follows from the proof of the $D$-equivalence conjecture for hyperkähler manifolds of $\operatorname{K3}^{[n]}$-type of [38]. Finally, if the cubic fourfolds satisfy Derived Torelli, we also have DF $\implies$ FM by [3, Corollary 9.3]. The question of whether implication FM $\implies$ BE holds is a conjecture by Huybrechts [24, Conjecture 7.3.21], and the implication BF $\implies$ BE is conjectured in [8, Conjecture 1.2]. For a survey on these equivalence relations, we refer to [8].
[…]
Finally, we prove that $\mathbb{L}$-equivalence classes of cubic fourfolds are finite:
Theorem 1.10 (Theorem 5.3).
For any cubic fourfold $X$, there exist finitely many cubic fourfolds $Y$ such that $X\sim_{\mathbb{L}}Y$.
Acknowledgements
We thank Franco Giovenzana for pointing out the paper [13] to us, and we would also like to thank Evgeny Shinder for telling us about Proposition 5.8. Finally, we thank Corey Brooke, Sarah Frei, and Lisa Marquand for useful comments on an earlier version of this paper. This work is a contribution to Project A22 in SFB 195 No. 286237555 of DFG. The second author is a member of the INdAM group GNSAGA.
2. Preliminaries
2.1. Lattice theory
Our main reference for lattice theory is [45]. A lattice is a free, finitely generated abelian group $L$ together with a non-degenerate symmetric integral bilinear form
$(-,-)\colon L\times L\to\mathbb{Z}.$
For elements $v,w\in L$, we often write $v^{2}\coloneqq(v,v)$ and $v\cdot w\coloneqq(v,w)$. A lattice $L$ is called even if $v^{2}$ is even for all $v\in L$.
Definition 2.1.
Let $v\in L$ be an element of a lattice. The divisibility of $v$ in $L$ is defined to be the positive integer
$\operatorname{div}(v)\coloneqq\gcd_{w\in L}(v\cdot w).$
The dual of $L$ is the group $L^{\vee}\coloneqq\operatorname{Hom}(L,\mathbb{Z})$, and we have a natural group homomorphism
$\begin{array}[]{rcl}L&\longrightarrow&L^{\vee}\\ v&\longmapsto&(v,-),\end{array}$
which is injective as $L$ is non-degenerate. The cokernel $A_{L}\coloneqq L^{\vee}/L$ is called the discriminant group. Via the embedding $L\hookrightarrow L\otimes\mathbb{Q}$, there is a natural identification
(2.1) $L^{\vee}=\left\{x\in L\otimes\mathbb{Q}\mid\forall v\in L:(x,v)\in\mathbb{Z}\right\}.$
This defines a natural symmetric bilinear form on $L^{\vee}$ taking values in $\mathbb{Q}$, which induces a bilinear form $b\colon A_{L}\times A_{L}\to\mathbb{Q}/\mathbb{Z}$. If $L$ is even, the bilinear form $b$ induces a quadratic form $q\colon A_{L}\to\mathbb{Q}/2\mathbb{Z}$. The pair $(A_{L},q)$ is called a finite quadratic module.
The order of $A_{L}$ is called the discriminant $\operatorname{disc}(L)$ of $L$. If $\operatorname{disc}(L)=1$, we say that $L$ is unimodular. If $M\in\operatorname{GL}_{n}(\mathbb{Z})$ is the matrix for the bilinear form on $L$ with respect to some $\mathbb{Z}$-basis of $L$, we have
$\operatorname{disc}(L)=|\det(M)|.$
If $L$ is a lattice, then $L\otimes\mathbb{Q}$ (resp. $L\otimes\mathbb{R}$) is a $\mathbb{Q}$- (resp. $\mathbb{R}$-)vector space which inherits a symmetric bilinear form from $L$. The signature $\operatorname{sgn}(L)=(\ell_{+},\ell_{-})$ is defined to be the signature of the real symmetric bilinear form on the $\mathbb{R}$-vector space $L\otimes\mathbb{R}$.
The genus $\mathcal{G}(N)$ of a lattice $N$ is the set of lattices $N^{\prime}$ satisfying each of the following:
$\operatorname{rk}(N)=\operatorname{rk}(N^{\prime}),\qquad\operatorname{sgn}(N)=\operatorname{sgn}(N^{\prime}),\qquad A_{N}\simeq A_{N^{\prime}}.$
For lattices $L$ and $N$, a group homomorphism $f\colon N\to L$ with the property that $(v,w)=(f(v),f(w))$ for all $v,w\in N$ is called a metric morphism. Bijective metric morphisms are called isometries. The group of isometries of $L$ is denoted $O(L)$. Similarly, we write $O(A_{L})$ for the group of isometries of the finite quadratic module $A_{L}$. Note that we have a natural group homomorphism
$\begin{array}[]{rcl}O(L)&\longrightarrow&O(A_{L})\\ f&\longmapsto&\overline{f}.\end{array}$
Metric morphisms are injective, since we assume our lattices to be non-degenerate. A metric morphism is called a primitive embedding if $L/N$ is a torsion-free abelian group. On the other hand, $f\colon N\to L$ is called an overlattice if $L/N$ is a finite abelian group. Two overlattices $f\colon N\hookrightarrow L$, $g\colon N^{\prime}\hookrightarrow L^{\prime}$ are called isomorphic if there exist isometries $\sigma\colon N\simeq N^{\prime}$, $\tau\colon L\simeq L^{\prime}$ fitting into a commutative square
[…]
Let $f\colon N\to L$ be an overlattice, and assume $N$ and $L$ are both even. The following results also have analogues for odd lattices [45]. From (2.1), we obtain a natural chain of finite-index injective group homomorphisms:
$N\hookrightarrow L\hookrightarrow L^{\vee}\hookrightarrow N^{\vee}.$
In particular, $G_{L}\coloneqq L/N$ can naturally be seen as a subgroup of $A_{N}$. Note that $L/N$ is an isotropic subgroup, and that $G_{L}^{\perp}=L^{\vee}/N$. This means that we have an isometry of finite quadratic modules $G_{L}^{\perp}/G_{L}\simeq A_{L}$.
Lemma 2.2.
[45, Proposition 1.4.2] Let $f\colon N\to L$ and $g\colon N\to L^{\prime}$ be two even overlattices of an even lattice $N$.
  1. (1)
    There is an isometry $\tau\colon L\simeq L^{\prime}$ fitting into a commutative triangle
[…]
    if and only if $G_{L}=G_{L^{\prime}}$ as subgroups of $A_{N}$.
  2. (2)
    Given an isometry $\sigma\colon N\simeq N$, there exists an isometry $\tau\colon L\simeq L^{\prime}$ fitting into a commutative square
[…]
    if and only if $\overline{\sigma}(G_{L})=G_{L^{\prime}}$.
  3. (3)
    The assignment $(f\colon N\hookrightarrow L)\mapsto G_{L}$ is a bijection between the set of isomorphism classes of overlattices of $N$, and the set of $O(N)$-orbits of isotropic subgroups of $A_{N}$.
Overlattices play a crucial role in the study of primitive embeddings of lattices, for the following reason. If $f\colon N\to L$ is a primitive embedding, let us denote its orthogonal complement by
$T=N^{\perp}\coloneqq\left\{v\in L\mid\forall n\in N:v\cdot n=0\right\}\subset L.$
Then $T\hookrightarrow L$ is also a primitive embedding. However, the metric morphism $N\oplus T\hookrightarrow L$ is an overlattice. By the above discussion, we obtain a natural chain of finite-index embeddings
$N\oplus T\hookrightarrow L\hookrightarrow L^{\vee}\hookrightarrow N^{\vee}\oplus T^{\vee}.$
Thus the finite group $G_{L}\coloneqq L/(N\oplus T)$ is naturally a subgroup of $A_{N}\oplus A_{T}$, and we call it the gluing group of the primitive embedding $f$. The natural projection maps $A_{N}\oplus A_{T}\to A_{N}$ and $A_{N}\oplus A_{T}\to A_{T}$ restrict to injective group homomorphisms $i_{N}\colon G_{L}\hookrightarrow A_{N}$ and $i_{T}\colon G_{L}\hookrightarrow A_{T}$, respectively.
To summarise, the primitive embedding $f\colon N\hookrightarrow L$ determines a pair $(G_{f},\gamma_{f}),$ where $G_{f}\coloneqq i_{N}(G_{L})\subset A_{N}$ is a subgroup, and $\gamma_{f}\coloneqq i_{T}\circ i_{N}^{-1}\colon G_{L}\to A_{T}$ is an injective group homomorphism, and the graph $\Gamma_{\gamma_{f}}\subset A_{N}\oplus A_{T}$ is precisely the gluing group $G_{L}$, which is an isotropic subgroup that satisfies $G_{L}^{\perp}/G_{L}\simeq A_{L}$.
Remark 2.3.
Conversely, suppose $(G,\gamma)$ is a pair consisting of a subgroup $G\subset A_{N}$ and an injective group homomorphism $\gamma\colon G\hookrightarrow A_{T}$ with the property that the graph $\Gamma_{\gamma}\subset A_{N}\oplus A_{T}$ is an isotropic subgroup which satisfies $\Gamma_{\gamma}^{\perp}/\Gamma_{\gamma}\simeq A_{L}$. Then we obtain an overlattice $N\oplus T\hookrightarrow L^{\prime}$ corresponding to $\Gamma_{\gamma}$ via Lemma 2.2. Moreover, the metric morphism $N\hookrightarrow N\oplus T\hookrightarrow L^{\prime}$ is a primitive embedding with orthogonal complement equal to the image of $T$ in $L^{\prime}$.
By construction, the lattice $L^{\prime}$ satisfies $A_{L^{\prime}}\simeq\Gamma_{\gamma}^{\perp}/\Gamma_{\gamma}\simeq A_{L}$. Unfortunately, $L^{\prime}$ is not necessarily isometric to $L$. However, since $L$ and $L^{\prime}$ have the same rank, signature and discriminant group, they are in the same genus.
Lemma 2.4.
[45, Proposition 1.5.1] Let $N$ be an even lattice and let $f\colon N\hookrightarrow L$ and $g\colon N\hookrightarrow L^{\prime}$ be primitive embeddings into even lattices $L$, $L^{\prime}$. Write $T\coloneqq N^{\perp}\subset L$ and $T^{\prime}\coloneqq N^{\perp}\subset L^{\prime}$.
  1. (1)
    There is an isometry $\tau\colon L\simeq L^{\prime}$ fitting into a commutative triangle
[…]
    if and only if there is an isometry $\sigma_{T}\colon T\simeq T^{\prime}$ such that $(G_{f},\overline{\sigma}_{T}\circ\gamma_{f})=(G_{g},\gamma_{g})$.
  2. (2)
    Fix an isometry $\sigma_{N}\colon N\simeq N$. There is an isometry $\tau\colon L\simeq L^{\prime}$ fitting into a commutative square
[…]
    if and only if there is an isometry $\sigma_{T}\colon T\simeq T^{\prime}$ such that the diagram
    ${G_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma_{f}}$$\scriptstyle{\overline{\sigma}_{N}}$${A_{T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{\sigma}_{T}}$${G_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma_{g}}$${A_{T^{\prime}}}$
    commutes.
The following lemma was noted in [21, §14.0], and was a crucial part of the technical computations in [41].
Lemma 2.5.
Let $f\colon N\hookrightarrow L$ be a primitive embedding with orthogonal complement $T=N^{\perp}\subset L$ and gluing group $G\subset A_{N}\oplus A_{T}$. Then we have
$\operatorname{disc}(T)\operatorname{disc}(N)=|G|^{2}\operatorname{disc}(L).$
Proof.
Recall that we have the following chain of embeddings
$N\oplus T\hookrightarrow L\hookrightarrow L^{\vee}\hookrightarrow N^{\vee}\oplus T^{\vee}.$
We have $L:N\oplus T]=|G|$ by construction, hence also $N^{\vee}\oplus T^{\vee}:L^{\vee}]=|G|$. The embedding $L\hookrightarrow L^{\vee}$ has index $\operatorname{disc}(L)$, by definition. In particular, the index of the composition of the three embeddings is $N^{\vee}\oplus T^{\vee}:N\oplus T]=|G|^{2}\operatorname{disc}(L)$. On the other hand, we have
$\frac{N^{\vee}\oplus T^{\vee}}{N\oplus T}\simeq A_{N\oplus T}\simeq A_{N}\oplus A_{T},$
hence $N^{\vee}\oplus T^{\vee}:N\oplus T]=\operatorname{disc}(N)\operatorname{disc}(T).$ To summarise, we have
$\operatorname{disc}(N)\operatorname{disc}(T)=[N^{\vee}\oplus T^{\vee}:N\oplus T]=|G|^{2}\operatorname{disc}(L),$
as required. ∎
If $L$ is a lattice of rank $1$, and $v\in L$ is a generator with $v^{2}=n\in\mathbb{Z}$, then we usually denote $L$ by $\langle n\rangle$. Another important example of a lattice is the root lattice $A_{2}$, which is the lattice of rank $2$ given by the intersection matrix
$A_{2}\coloneqq\left(\begin{matrix}2&-1\\ -1&2\end{matrix}\right).$
Finally, the hyperbolic plane, denoted $U$, is the indefinite even unimodular lattice of rank 2. Its intersection matrix is
$U\coloneqq\left(\begin{matrix}0&1\\ 1&0\end{matrix}\right).$
Remark 2.6.
Let $q\colon C_{3}\to\mathbb{Q}/2\mathbb{Z}$ be a non-degenerate finite quadratic form of order $3$. Then if $x\in C_{3}$ is a generator, we see that $q(x)=\frac{a}{3}\pmod{2\mathbb{Z}}$, where $a\in(\mathbb{Z}/3\mathbb{Z})^{\times}$. Therefore, $q(2x)=\frac{4a}{3}\equiv\frac{a}{3}\pmod{2\mathbb{Z}}$. As a consequence, $C_{3}$ and $C_{3}(-1)$ are the only two isomorphism classes of non-degenerate finite quadratic forms of order $3$.
We now compute the signatures of these finite quadratic forms. Suppose $C_{3}$ is the finite quadratic module of order $3$ with $q(x)=\frac{2}{3}\pmod{2\mathbb{Z}}$ for $x\in C_{3}$ a generator. We can find an explicit lattice whose discriminant group is $C_{3}$. Indeed, the lattice $A_{2}$ is a positive-definite lattice of rank $2$ and discriminant $3$, and we can compute that, for a generator $y\in A_{A_{2}}$, we have $q(y)=\frac{2}{3}\pmod{2\mathbb{Z}}$. In other words, the discriminant group of $A_{2}$ is precisely $C_{3}$. Therefore, we have
$\operatorname{sgn}(C_{3})\equiv\operatorname{sgn}(A_{2})\equiv 2\pmod{8}$,
and also $\operatorname{sgn}(C_{3}(-1))\equiv\operatorname{sgn}(A_{2}(-1))\equiv-2\equiv 6\pmod{8}$.
2.2. Cubic fourfolds and K3 surfaces
Our main references for K3 surfaces and cubic fourfolds are [21, 24]. Let $X$ be a cubic fourfold. We consider the cohomology group $H^{4}(X,\mathbb{Z})$. The intersection product defines a lattice structure on $H^{4}(X,\mathbb{Z})$, which abstractly makes it isometric to the lattice
$\Lambda_{\text{cub}}\coloneqq E_{8}^{\oplus 2}\oplus U^{\oplus 2}\oplus\langle 1\rangle^{\oplus 3}.$
Here, $E_{8}$ denotes the unique positive-definite unimodular lattice of rank $8$. The signature of $\Lambda_{\text{cub}}$ is $(21,2)$. We denote $h^{2}\coloneqq(1,1,1)\in\langle 1\rangle^{\oplus 3}$. Then we define
$\Lambda_{\text{cub}}^{0}\coloneqq(h^{2})^{\perp}\simeq E_{8}^{\oplus 2}\oplus U^{\oplus 2}\oplus A_{2}.$
If $H^{2}\in H^{4}(X,\mathbb{Z})$ is the square of a hyperplane, then the primitive cohomology $H^{4}(X,\mathbb{Z})_{\text{prim}}\coloneqq(H^{2})^{\perp}\subset H^{4}(X,\mathbb{Z})$ is isometric to $\Lambda_{\text{cub}}^{0}$ as a lattice. Note that $\Lambda_{\text{cub}}$ is not an even lattice, but $\Lambda_{\text{cub}}^{0}$ is.
The Hodge structure on $H^{4}(X,\mathbb{Z})$ is of K3-type. That is,
$\begin{array}[]{lll}H^{4,0}(X)&\simeq H^{0,4}(X)&=0\\ H^{3,1}(X)&\simeq H^{1,3}(X)&\simeq\mathbb{C}.\end{array}$
Moreover, since $H^{2}$ is of type $(2,2)$, the primitive cohomology $H^{4}(X,\mathbb{Z})_{\text{prim}}$ inherits a Hodge structure from $H^{4}(X,\mathbb{Z})$ which is also of K3-type.
Definition 2.7.
A lattice with a Hodge structure is called a Hodge lattice. If $H_{1},H_{2}$ are Hodge lattices, then a Hodge isometry is an isomorphism of groups $f\colon H_{1}\simeq H_{2}$ which is simultaneously an isomorphism of Hodge structures and an isometry.
The algebraic part of $H^{4}(X,\mathbb{Z})$ is denoted
$A(X)\coloneqq H^{2,2}(X)\cap H^{4}(X,\mathbb{Z}).$
This is a positive-definite lattice, which contains the class $H^{2}$. Since the self-intersection number of $H^{2}$ is $3$, it follows that $A(X)$ is not an even lattice. Finally, we denote
$A(X)_{\text{prim}}\coloneqq(H^{2})^{\perp}\subset A(X).$
As opposed to the full algebraic lattice, the primitive part $A(X)_{\text{prim}}$ is an even lattice, since it is a sublattice of the primitive cohomology group $H^{4}(X,\mathbb{Z})_{\mathrm{prim}}$, which is even.
Definition 2.8.
The transcendental lattice of a cubic fourfold is the Hodge lattice $T(X)\coloneqq A(X)^{\perp}\subset H^{4}(X,\mathbb{Z})$. Equivalently, we may define $T(X)=A(X)_{\text{prim}}^{\perp}\subset H^{4}(X,\mathbb{Z})_{\text{prim}}$.
Theorem 2.9 (Torelli Theorem for cubic fourfolds).
[54] Let $X$ and $Y$ be two cubic fourfolds. Then $X$ and $Y$ are isomorphic if and only if there exists a Hodge isometry $H^{4}(X,\mathbb{Z})_{\text{prim}}\simeq H^{4}(Y,\mathbb{Z})_{\text{prim}}$.
A remarkable class of divisors in the moduli space of cubic fourfolds are the Noether–Lefschetz divisors $\mathcal{C}_{d}$, consisting of so called special cubic fourfolds. A cubic fourfold $X$ is called special if its algebraic lattice $A(X)$ has rank at least $2$, or equivalently if $X$ contains a surface which is not homologous to a complete intersection. As a consequence we have a primitive saturated lattice of rank $2$ containing the square of the hyperplane class, and the class $\alpha$ of this surface:
$K=\langle h^{2},\alpha\rangle\subset A(X).$
The locus of special cubic fourfolds with $\operatorname{disc}(K)=d$ is denoted by $\mathcal{C}_{d}$.
A fundamental tool for studying cubic fourfolds is the period map, which associates a cubic fourfold $X$ in its moduli space to its primitive cohomology $H^{4}_{\operatorname{prim}}(X,\mathbb{Z})$, viewed as a point in the corresponding period domain, see [24] for more details.
Theorem 2.10 (Surjectivity of the period map for cubic fourfolds).
[34, 33] The image of the period map for cubic fourfolds is the complement of the divisors $\mathcal{C}_{2}\cup\mathcal{C}_{6}$.
It follows from the surjectivity of the period map that $A(X)_{\mathrm{prim}}$ contains no vectors of square $2$, nor any vectors of square $6$ and divisibility $3$ in $H^{4}(X,\mathbb{Z})_{\mathrm{prim}}$. For this reason, we define:
Definition 2.11.
Let $N$ be a lattice. We denote
$\mathcal{H}(N)\coloneqq\left\{N^{\prime}\in\mathcal{G}(N)\mid\nexists v\in N^{\prime}:v^{2}=2\text{ or }v^{2}=6\wedge\operatorname{div}(v)=3\right\}.$
The derived category of a cubic fourfold $X$ admits a semiorthogonal decomposition [32]:
$\mathcal{D}^{b}(X)\simeq\langle\mathcal{A}_{X},\mathcal{O}_{X},\mathcal{O}_{X}(1),\mathcal{O}_{X}(2)\rangle.$
The component $\mathcal{A}_{X}$ is a K3 category called the Kuznetsov component. The topological Grothendieck group $K_{\text{top}}(\mathcal{A}_{X})$ has a natural Hodge structure of K3-type of weight $2$ as defined in [2]. The resulting Hodge lattice is usually denoted $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$. As a lattice, $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is isometric to the even, unimodular lattice of signature $(20,4)$:
$\tilde{\Lambda}\coloneqq U^{\oplus 4}\oplus E_{8}^{\oplus 2}.$
The transcendental lattice $T(X)$ has a natural embedding into $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$. In fact, $T(X)$ is the transcendental sublattice of $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ which is a notion that we will define in Section 2.3.
We now turn our attention to K3 surfaces. We assume all our K3 surfaces to be projective. If $S$ is a K3 surface, the intersection pairing and Hodge structure turn the cohomology group $H^{2}(S,\mathbb{Z})$ into a Hodge lattice of K3-type. The underlying lattice structure of this Hodge lattice is
$\Lambda_{\operatorname{K3}}\coloneqq U^{\oplus 3}\oplus E_{8}(-1)^{\oplus 2}.$
This is an even unimodular lattice of signature $(3,19)$. The algebraic part of $H^{2}(S,\mathbb{Z})$ is called the Néron–Severi lattice of $S$, denoted
$\operatorname{NS}(S)\coloneqq H^{1,1}(S)\cap H^{2}(S,\mathbb{Z}).$
The Néron–Severi lattice has signature $(1,\rho-1)$, where $\rho=\operatorname{rk}\operatorname{NS}(S)$ is the Picard rank of $S$. Its orthogonal complement is the transcendental lattice
$T(S)\coloneqq\operatorname{NS}(S)^{\perp}\subset H^{2}(S,\mathbb{Z}).$
Similarly to cubic fourfolds, the Torelli Theorem for K3 surfaces states that two K3 surfaces $S$ and $S^{\prime}$ are isomorphic if and only if the cohomology groups $H^{2}(S,\mathbb{Z})$ and $H^{2}(S^{\prime},\mathbb{Z})$ are Hodge isometric.
For a K3 surface $S$, we consider its full cohomology group $H^{*}(S,\mathbb{Z})\simeq H^{0}(S,\mathbb{Z})\oplus H^{2}(S,\mathbb{Z})\oplus H^{4}(S,\mathbb{Z})$. Via the natural isomorphisms $H^{0}(S,\mathbb{Z})\simeq\mathbb{Z}$ and $H^{4}(S,\mathbb{Z})\simeq\mathbb{Z}$, we usually write an element of $H^{*}(S,\mathbb{Z})$ as $(r,\ell,s)$, where $r,s\in\mathbb{Z}$ and $\ell\in H^{2}(S,\mathbb{Z})$. We can put a lattice structure on $H^{*}(S,\mathbb{Z})$ by the following definition:
$(r,\ell,s)\cdot(r^{\prime},\ell^{\prime},s^{\prime})=-2(rs^{\prime}+r^{\prime}s)+\ell\cdot\ell^{\prime}.$
In other words, we declare that $H^{0}(S,\mathbb{Z})\oplus H^{4}(S,\mathbb{Z})$ gets the structure of a hyperbolic plane $U$ which is orthogonal to $H^{2}(S,\mathbb{Z})$. There is also a natural Hodge structure on $H^{*}(S,\mathbb{Z})$ which it inherits from $H^{2}(S,\mathbb{Z})$. With this Hodge structure and lattice structure, we denote this Hodge lattice by $\tilde{H}(S,\mathbb{Z})$ and call it the Mukai lattice of $S$.
We recall the Derived Torelli Theorem by Mukai and Orlov, which is the most fundamental result regarding derived equivalence for K3 surfaces:
Theorem 2.12 (Derived Torelli Theorem for K3 surfaces).
[43, 48] Let $S$ and $S^{\prime}$ be K3 surfaces. The following are equivalent:
  1. (1)
    $S$ and $S^{\prime}$ are Fourier–Mukai partners, that is, there is an equivalence $\mathcal{D}^{b}(S)\simeq\mathcal{D}^{b}(S^{\prime})$.
  2. (2)
    There is a Hodge isometry $T(S)\simeq T(S^{\prime})$.
  3. (3)
    There is a Hodge isometry $\tilde{H}(S,\mathbb{Z})\simeq\tilde{H}(S^{\prime},\mathbb{Z})$.
Returning to cubic fourfolds, it is known for several families of special cubic fourfolds that the Kuznetsov component is explicitly realised as the derived category of a (possibly twisted) K3 surface: for cubic fourfolds containing a plane this goes back to [32] and has been extended to the non-generic nodal case in [42]. Recently this has been studied for cubic fourfolds containing multiple planes in [17]. However, it is currently an open conjecture that Theorem 2.12 has an analogue for the Kuznetsov components of cubic fourfolds. That is, for two cubic fourfolds $X,Y$, it is currently not known whether there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ if there is a Hodge isometry $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})\simeq\tilde{H}(\mathcal{A}_{Y},\mathbb{Z})$. There is the following positive result, however:
Theorem 2.13.
[20, 2] Let $X$ and $Y$ be cubic fourfolds. Consider the following three statements:
  1. (1)
    There is a Fourier–Mukai equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$.
  2. (2)
    There is a Hodge isometry $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})\simeq\tilde{H}(\mathcal{A}_{Y},\mathbb{Z})$.
  3. (3)
    There is a Hodge isometry $T(X)\simeq T(Y)$.
Then we have
$(1)\implies(2)\implies(3).$
Moreover, suppose that there is a twisted K3 surface $(S,\alpha)$ and a Hodge isometry $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})(-1)\simeq\tilde{H}(S,\alpha,\mathbb{Z})$. Then in addition, we have $(2)\implies(1).$
2.3. Hodge realisation
We briefly recall the constructions of [11, 40, 41]. We denote by $\operatorname{HS}_{\mathbb{Z}}$ the category of pure, graded, polarisable, integral Hodge structures. The Grothendieck group of Hodge structures is defined to be the group
$K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z}})\coloneqq\mathbb{Z}[\operatorname{HS}_{\mathbb{Z}}/_{\simeq}]\big/([H_{1}\oplus H_{2}]-[H_{1}]-[H_{2}]).$
There is a Hodge realisation map
$\operatorname{Hdg}_{\mathbb{Z}}\colon K_{0}(\operatorname{Var}_{\mathbb{C}})\to K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z}})$
defined by sending the class of a smooth, projective variety $[X]$ to the class of its full cohomology group $H^{\bullet}(X,\mathbb{Z})]$, considered with its Hodge structure.
Remark 2.14.
The Hodge realisation map factors through $K_{0}(\operatorname{Var}_{\mathbb{C}})[\mathbb{L}^{-1}]$. In other words, if $X$ and $Y$ are smooth, projective, $\mathbb{L}$-equivalent varieties, then we have $H^{\bullet}(X,\mathbb{Z})]=[H^{\bullet}(Y,\mathbb{Z})]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z}})$.
For any $n\in\mathbb{Z}$, we have a full subcategory $\operatorname{HS}_{\mathbb{Z},n}$ of $\operatorname{HS}_{\mathbb{Z}}$, consisting of pure, polarisable, integral Hodge structures of weight $n$. This induces group homomorphisms for each $n$:
$\begin{array}[]{rcl}K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z}})&\longrightarrow&K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},n})\\ \text{[}H^{\bullet}]&\longmapsto&[H^{n}].\end{array}$
Suppose we have $H\in\operatorname{HS}_{\mathbb{Z},2n}$ a Hodge structure of weight $2n$, then the Tate twist H(n1)H(n-1) is a Hodge structure of weight $2$. This yields an isomorphism
$K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2n})\simeq K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2}).$
For a Hodge structure $H\in\operatorname{HS}_{\mathbb{Z},2}$ of weight $2$, we denote by $T(H)$ the transcendental sub-Hodge structure of $H$. That is, $T(H)$ is the minimal, integral, primitive sub-Hodge structure with the property that $H^{2,0}\subset T(H)_{\mathbb{C}}$. Moreover, we write
$\operatorname{NS}(H)\coloneqq H^{1,1}\cap H.$
Example 2.15.
If $X$ is a hyperkähler manifold, then
$\operatorname{NS}(H^{2}(X,\mathbb{Z}))=\operatorname{NS}(X),\qquad T(H^{2}(X,\mathbb{Z}))=T(X).$
If $Y$ is a cubic fourfold, then $\operatorname{NS}(H^{4}(Y,\mathbb{Z})(1))=A(Y)$ and $T(H^{4}(Y,\mathbb{Z})(1))=T(Y)(1)$.
Definition 2.16.
For $H\in\operatorname{HS}_{\mathbb{Z},2}$, we call the group
$G(H)\coloneqq\frac{H}{\operatorname{NS}(H)\oplus T(H)}$
the gluing group of $H$.
Lemma 2.17.
[40, 41] There is a group homomorphism
$\begin{array}[]{rcl}K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2})&\longrightarrow&\mathbb{Q}^{\times}\\ {[}H]&\longmapsto&\left|G(H)_{\text{tors}}\right|.\end{array}$
Lemma 2.18.
If $X$ is a cubic fourfold, then we have
$|G(H^{4}(X,\mathbb{Z})(1))|=\operatorname{disc}(T(X)).$
Proof.
The transcendental lattice $T(X)$ is a primitive sublattice of the unimodular lattice $H^{4}(X,\mathbb{Z})$, hence the gluing group is isomorphic to $A_{T(X)}$ by Remark 2.3. ∎
Lemma 2.19.
Let $F(X)$ be the Fano variety of lines on a cubic fourfold $X$. Then we have
$|G(F(X))|=\operatorname{disc}(T(F(X)))=\operatorname{disc}(T(X)).$
Proof.
Addington showed in [1, Corollary 8] that there is a natural embedding $A_{2}=\langle\lambda_{1},\lambda_{2}\rangle\subset\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ such that
(2.2) $H^{2}(F(X),\mathbb{Z})\simeq(\lambda_{1}^{\perp})(-1)\subset\tilde{H}(\mathcal{A}_{X},\mathbb{Z})(-1),$
where the (1)(-1) indicates that we rescaled the bilinear forms by 1-1. Since the divisibility of $\lambda_{1}$ in $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$ is $1$, it follows that the coarseness of $F(X)$ is $\operatorname{crs}(F(X))=1$ (see [41, Definition 2.4] for the definition of coarseness). Therefore, by [41, Theorem 3.5], we find
$|G(F(X))|=\frac{\operatorname{disc}(T(F(X)))}{\operatorname{crs}(F(X))}=\operatorname{disc}(T(F(X))),$
as required. Moreover, by (2.2), $T(F(X))$ is Hodge anti-isometric to the transcendental sublattice of $\tilde{H}(\mathcal{A}_{X},\mathbb{Z})$, which is $T(X)$, hence $\operatorname{disc}(T(F(X)))=\operatorname{disc}(T(X))$. ∎
Suppose that there is a K3 surface $S$ with a Hodge isometry $T(X)(-1)\simeq T(S)$, so that $F(X)$ is a moduli space of stable sheaves on $S$ by [3, Corollary 9.6]. Then Lemma 2.19 is equivalent to saying that $F(X)$ is a fine moduli space of sheaves on $S$ by [39, Theorem 3.9].
The following result appears as Lemma 3.2 in [4].
Theorem 2.20.
Let $X$ and $Y$ be cubic fourfolds (resp. K3 surfaces). Suppose $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $[T(X)]=[T(Y)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z}})$, there is a Hodge isometry $T(X)\simeq T(Y)$.
Proof.
If $X$ and $Y$ are K3 surfaces, the statement is follows from [40, Proposition 3.11] combined with [11, Lemma 3.7].
If $X$ and $Y$ are cubic fourfolds, we have $[T(X)(1)]=[T(Y)(1)]$ in $K_{0}^{\oplus}(\operatorname{HS}_{\mathbb{Z},2})$. Since $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$, this implies by [11, Lemma 3.7] that there exists an isomorphism of Hodge structures (not necessarily a Hodge isometry) $T(X)(1)\simeq T(Y)(1)$. Moreover, since $X$ and $Y$ are $\mathbb{L}$-equivalent, we have $\operatorname{disc}(T(X))=\operatorname{disc}(T(Y))$ by Lemma 2.17 and Lemma 2.18. It now follows from [40, Corollary 3.10] that we have a Hodge isometry $T(X)(1)\simeq T(Y)(1)$, which clearly induces a Hodge isometry $T(X)\simeq T(Y)$ via the Tate twist. ∎
3. Counting Fourier–Mukai partners of cubic fourfolds
We compute a counting formula for Fourier–Mukai partners of cubic fourfolds with certain properties. The results in this section are closely related to [19, 35, 51, 13, 9].
We begin by collecting a few remarks and criteria concerning cubic fourfolds that satisfy Derived Torelli in the sense of Definition 1.7, and we describe how to find an example of a cubic which doesn’t.
Remark 3.1.
Notice that the implication $(1)\Rightarrow(2)$ in Definition 1.7 always holds. By [23, Proposition 3.4], any equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ induces an orientation-preserving Hodge isometry $\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})\to\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})$ ([23, Theorem 1.2]). Then, by [2, Proposition 2.3], there is an identification $H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cong A_{2}^{\perp}\subset\widetilde{H}(\mathcal{A}_{X},\mathbb{Z})$. By Voisin’s proof of the integral Hodge conjecture for cubic fourfolds, the primitive algebraic lattices coincide with the primitive Hodge classes, $A(X)_{\mathrm{prim}}=H^{4}_{\mathrm{prim}}(X,\mathbb{Z})\cap H^{2,2}(X)$. Hence we have a Hodge isometry $T(X)=A(X)_{\mathrm{prim}}^{\perp}\cong A(Y)_{\mathrm{prim}}^{\perp}=T(Y)$.
The following result follows immediately from Theorem 2.13.
Proposition 3.2.
In each of the following cases, the cubic fourfold $X$ satisfies Derived Torelli:
  1. (1)
    There is a twisted K3 surface $(S,\alpha)$ such that there exists a Hodge isometry $\tilde{H}(S,\alpha,\mathbb{Z})\simeq\tilde{H}(\mathcal{A}_{X},\mathbb{Z})(-1)$, and there is a unique embedding $T(X)\hookrightarrow\tilde{\Lambda}$.
  2. (2)
    If there exists a K3 surface $S$ and a Hodge isometry $T(X)\simeq T(S)(-1)$.
Note that (2) is a special case of (1) in Proposition 3.2.
Remark 3.3.
Notice that a weaker Derived Torelli statement can be found in Conjecture 6.1 of [25], where the existence of an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ is conjectured to be equivalent to an orientation preserving Hodge isometry $\widetilde{H}(X,\mathbb{Z})\simeq\widetilde{H}(Y,\mathbb{Z})$. As Hodge isometries of $T(X)$ need not to lift to Hodge isometries of the Mukai lattice, it is reasonable to expect that some cubic fourfolds do not satisfy Derived Torelli in the sense of Definition 1.7.
Remark 3.4.
Let $T$ be a lattice such that $\operatorname{disc}(T)$ is not divisible by $3$, suppose that we have a primitive embedding $T\hookrightarrow\Lambda_{\text{cub}},$ and let $N\coloneqq T^{\perp}\subset\Lambda_{\text{cub}}$. Let $G=\frac{\Lambda}{T\oplus N}$ be the gluing group for this primitive embedding. Then, by Lemma 2.5, we have
$\frac{\operatorname{disc}(T)}{|G|}\cdot\frac{\operatorname{disc}(N)}{|G|}=\operatorname{disc}(\Lambda_{\text{cub}}^{0})=3.$
Since $\operatorname{disc}(T)$ is not divisible by $3$, and $|G|$ divides both $\operatorname{disc}(T)$ and $\operatorname{disc}(N)$, it follows that we have
$\operatorname{disc}(T)=|G|,\qquad\operatorname{disc}(N)=3\cdot|G|.$
In particular, the canonical embedding $G\hookrightarrow A_{T}$ is actually a group isomorphism $G\simeq A_{T}$. Therefore, the natural embedding $G\hookrightarrow A_{N}$ induces an isometry $A_{N}\simeq A_{T}(-1)\oplus C_{3}$, where $C_{3}$ is one of the two non-degenerate quadratic modules of order $3$ (see Remark 2.6). As a consequence, we have a canonical isomorphism
(3.1) $O(A_{N})\simeq O\left(A_{T}(-1)\right)\oplus O(C_{3})\simeq O\left(A_{T}\right)\oplus\left\{\pm\operatorname{id}_{C_{3}}\right\}.$
From this observation, we obtain an action of $O(N)$ on $O(A_{T})$. Indeed, if $f\in O(N)$ is an isometry, it induces an isometry $\overline{f}\in O(A_{N})$, which can be uniquely written as $\overline{f}=(\sigma,\tau)\in O(A_{T})\oplus O(C_{3})$ via (3.1).
On the other hand, if $\operatorname{disc}(N)$ is not divisible by $3$, by the symmetry of the situation we obtain an action of $O(T)$ on $O(A_{N})$.
The main result in this section is the following counting formula:
Theorem 3.5.
Let $X$ be a cubic fourfold that satisfies Derived Torelli (See Definition 1.7). Assume moreover that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
[…]
Here, $N=A(X)_{\text{prim}}$ and the action of $O(N^{\prime})$ on $O(A_{T(X)})$ is the one defined by Remark 3.4.
Remark 3.6.
If one replaces $\mathcal{H}(N)$ by $\mathcal{G}(N)$ in (3.2), one obtains the number of virtual Fourier–Mukai partners as in [9].
Definition 3.7.
For lattices $T,N,L$, we denote by $\operatorname{Emb}_{N}(T,L)$ the set of $O(\Lambda_{\text{cub}}^{0})$-orbits of primitive embeddings $f\colon T\hookrightarrow L$ such that $T^{\perp}\simeq N$. Note that $O(T)$ acts naturally on $\operatorname{Emb}_{N}(T,L)$.
We now prepare for the proof of Theorem 3.5 with some technical results.
Lemma 3.8.
Let $X$ be a cubic fourfold such that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then, for any Fourier–Mukai partner $Y$ of $X$, we have
$A(Y)_{\text{prim}}\in\mathcal{H}\left(A(X)_{\text{prim}}\right).$
In particular, $A(Y)_{\text{prim}}$ is in the same genus as $A(X)_{\text{prim}}$.
Proof.
To ease notation, we write $N=A(X)_{\text{prim}}$ and $N^{\prime}=A(Y)_{\text{prim}}$. By Remark 3.1, since $Y$ is a Fourier–Mukai partner of $X$, there exists a Hodge isometry $T(X)\simeq T(Y)$. In particular, just as in Remark 3.4, we find $A_{N}\simeq A_{T(X)}(-1)\oplus C_{3}$, where $C_{3}$ is one of the two non-degenerate quadratic modules of order $3$. Likewise, we obtain $A_{N^{\prime}}\simeq A_{T(Y)}(-1)\oplus C_{3}^{\prime},$ where $C_{3}^{\prime}$ is one of the two non-degenerate quadratic modules of order $3$. By Remark 2.6, we see that $C_{3}\simeq C_{3}^{\prime}$ or $C_{3}\simeq C_{3}^{\prime}(-1)$. We now argue that we must have $C_{3}\simeq C_{3}^{\prime}$, which implies $A_{N}\simeq A_{N^{\prime}}$, thus $N^{\prime}$ is in the genus of $N$. By construction, we have $\operatorname{sgn}(N)=\operatorname{sgn}(N^{\prime})$, thus $\operatorname{sgn}(A_{N})\equiv\operatorname{sgn}(A_{N^{\prime}})\pmod{8}$. This implies that we have
$0=\operatorname{sgn}(A_{N})-\operatorname{sgn}(A_{N^{\prime}})=\operatorname{sgn}(A_{T(X)(-1)})+\operatorname{sgn}(C_{3})-\operatorname{sgn}(A_{T(Y)(-1)})-\operatorname{sgn}(C_{3}^{\prime}).$
Since $T(X)$ is isometric to $T(Y)$, we have $\operatorname{sgn}(A_{T(X)(-1)})=\operatorname{sgn}(A_{T(Y)(-1)})$, hence
$\operatorname{sgn}(C_{3})=\operatorname{sgn}(C_{3}^{\prime}).$
Recall from Remark 2.6 that the signature of $C_{3}$ is either $2$ or $6$ modulo $8$. In particular, we find $\operatorname{sgn}(C_{3}(-1))=-\operatorname{sgn}(C_{3})\not\equiv\operatorname{sgn}(C_{3})\pmod{8}$. Since $\operatorname{sgn}(C_{3})=\operatorname{sgn}(C_{3}^{\prime})$, we obtain $C_{3}\simeq C_{3}^{\prime}$, which shows that $N$ and $N^{\prime}$ are in the same genus.
We now check that $N^{\prime}\in\mathcal{H}(N)$. Firstly, note that if $N^{\prime}$ were to contain a vector of square $2$, this would be a contradiction with the surjectivity of the period map, see Theorem 2.10.
Similarly, we now prove that $N^{\prime}$ cannot contain a vector of square $6$ and divisibility $3$. For the sake of contradiction, let us suppose that there exists a vector $v\in N^{\prime}$ with $v^{2}=6$ and $\operatorname{div}_{N^{\prime}}(v)=3$. Then the element $\frac{v}{3}\in A_{N^{\prime}}\simeq A_{T(Y)(-1)}\oplus C_{3}$ has order $3$, hence it is a generator for $C_{3}$ and, importantly, is not contained in $A_{T(Y)(-1)}$. The gluing group $G\coloneqq\frac{H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}}{N\oplus T(Y)}$ is equal to $A_{T(Y)(-1)}\oplus A_{T(Y)}$ as a subgroup of $A_{N^{\prime}}\oplus A_{T(Y)}$. Therefore, we have a natural isomorphism $A_{H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}}\simeq G^{\perp}/G\simeq C_{3}$, showing that $A_{H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}}$ is also generated by $\frac{v}{3}$. This implies that $\frac{v}{3}\in H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}^{\vee}$, hence $v$ must have divisibility $3$ in $H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}$. This contradicts the surjectivity of the period map, c.f. Theorem 2.10.
This shows that $N^{\prime}\in\mathcal{H}(N)$, as required. ∎
Definition 3.9.
For a cubic fourfold $X$ and a lattice $N$, we write $\operatorname{FM}(X,N)$ for the set of isomorphism classes of Fourier–Mukai partners $Y$ of $X$ such that there exists an isometry $A(Y)_{\text{prim}}\simeq N$.
Proposition 3.10.
Let $X$ be a cubic fourfold which satisfies Derived Torelli. Let $N^{\prime}$ be a lattice. Then there is a one-to-one correspondence between the set $\operatorname{FM}(X,N^{\prime})$ and the set of $O_{\mathrm{Hodge}}(T(X))$-orbits of primitive embeddings in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$.
Proof.
If $Y\in\operatorname{FM}(X,N^{\prime})$, and we fix any Hodge isometry $\psi\colon T(X)\simeq T(Y)$, and any isometry $\phi\colon H^{4}(Y,\mathbb{Z})_{\text{prim}}\simeq\Lambda_{\text{cub}}^{0}$ then the composition
$i\colon T(X)\overset{\psi}{\simeq}T(Y)\hookrightarrow H^{4}(Y,\mathbb{Z})_{\text{prim}}\overset{\phi}{\simeq}\Lambda_{\text{cub}}^{0}$
is a primitive embedding with $T(X)^{\perp}\simeq N^{\prime}$. However, $i$ depends on the choice of a Hodge isometry $\psi$ and the choice of an isometry $\phi\colon H^{4}(Y,\mathbb{Z})_{\text{prim}}\simeq\Lambda_{\text{cub}}^{0}$. However, if $\psi^{\prime}\colon T(X)\simeq T(Y)$ is another Hodge isometry, and $\phi^{\prime}\colon H^{4}(Y,\mathbb{Z})_{\text{prim}}\simeq\Lambda_{\text{cub}}^{0}$ is another isometry, and we denote by $i^{\prime}\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ the resulting primitive embedding, then the commutative diagram
${i\colon}$${T(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\psi}$$\scriptstyle{(\psi^{\prime})^{-1}\circ\psi}$${T(Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{4}(Y,\mathbb{Z})_{\text{prim}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\phi}$${\Lambda_{\text{cub}}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\phi^{\prime}\circ\phi^{-1}}$${i^{\prime}\colon}$${T(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\psi^{\prime}}$${T(Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{4}(Y,\mathbb{Z})_{\text{prim}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\phi^{\prime}}$${\Lambda_{\text{cub}}^{0}}$
shows that the primitive embeddings $i$ and $i^{\prime}$ are in the same $O_{\mathrm{Hodge}}(T(X))$-orbit. We will denote an embedding in this orbit by $i_{Y}$ from now on. This process thus yields a well-defined map:
$\begin{array}[]{lrcl}\operatorname{emb}:&\operatorname{FM}(X,N^{\prime})&\longrightarrow&\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})\big/O_{\mathrm{Hodge}}(T(X))\\ &Y&\longmapsto&i_{Y}\end{array}$
The injectivity of this map follows essentially from the Torelli Theorem: if $Y,Y^{\prime}\in\operatorname{FM}(X,N^{\prime})$, such that $i_{Y}$ and $i_{Y^{\prime}}$ are in the same $O_{\mathrm{Hodge}}(T(X))$-orbit, then there is a Hodge isometry $\psi\colon T(X)\simeq T(X)$ fitting into a commutative diagram:
${i_{Y}\colon}$${T(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\psi}$${T(Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{4}(Y,\mathbb{Z})_{\text{prim}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\phi}$${\Lambda_{\text{cub}}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${i_{Y^{\prime}}\colon}$${T(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${T(Y^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{4}(Y^{\prime},\mathbb{Z})_{\text{prim}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\phi^{\prime}}$${\Lambda_{\text{cub}}^{0}}$
Now $(\phi^{\prime})^{-1}\circ\phi\colon H^{4}(Y,\mathbb{Z})_{\text{prim}}\simeq H^{4}(Y^{\prime},\mathbb{Z})_{\text{prim}}$ is a Hodge isometry, so that $Y\simeq Y^{\prime}$ by the Torelli Theorem.
For surjectivity, note that any primitive embedding $i\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ induces a Hodge structure on $\Lambda_{\text{cub}}^{0}$. We denote the resulting Hodge lattice $L_{i}$, i.e. $L_{i}$ is the Hodge lattice whose underlying lattice structure is $\Lambda_{\text{cub}}^{0}$ and whose Hodge structure is induced by $i$. Since $T(X)^{\perp}\simeq N^{\prime}\in\mathcal{H}(N)$, it follows that $N^{\prime}$ contains no vectors of square $6$ and divisibility $3$, hence there is a unique cubic fourfold $Y$ for which there is a Hodge isometry $H^{4}(Y,\mathbb{Z})_{\text{prim}}\simeq L_{i}$. By construction, we have $T(Y)\simeq T(L_{i}(1))\simeq T(X)$, and $A(Y)_{\text{prim}}\simeq\operatorname{NS}(L_{i}(1))\simeq N^{\prime}$, where $T(L_{i}(1))$ and $\operatorname{NS}(L_{i}(1))$ are defined as in Section 2.3. Since $X$ satisfies Derived Torelli, it follows that $Y\in\operatorname{FM}(X,N^{\prime})$. Clearly, we have $i=\operatorname{emb}(Y)$. Therefore, the map $\operatorname{emb}$ is surjective, as required. ∎
Corollary 3.11.
Let $X$ be a cubic fourfold which satisfies Derived Torelli, and such that $\operatorname{disc}(T(X))$ is not divisible by $3$. Then we have
$\operatorname{FM}(X)=\bigsqcup_{N^{\prime}\in\mathcal{H}(N)}\operatorname{FM}(X,N^{\prime})\simeq\bigsqcup_{N^{\prime}\in\mathcal{H}(N)}\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})/O_{\mathrm{Hodge}}(T(X)).$
We are now ready to prove Theorem 3.5.
Proof of Theorem 3.5.
Let $N^{\prime}\in\mathcal{H}(N)$. Fix, once and for all, an isometry $A_{N^{\prime}}\simeq A_{T(X)}\oplus C_{3}$. Via this isometry, we will view $A_{T(X)}$ as a quadratic submodule of $A_{N^{\prime}}$.
Suppose $i\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ is a primitive embedding with orthogonal complement $T(X)^{\perp}\simeq N^{\prime}$. From Lemma 2.4, it follows that $i$ induces a pair $(G_{i},\gamma_{i})$, where $G_{i}\subset A_{T(X)}$ and $\gamma_{i}\colon G\hookrightarrow A_{N^{\prime}}$. Here, $G_{i}$ is isomorphic, as a group, to the gluing group
$G_{i}\simeq\frac{\Lambda_{\text{cub}}^{0}}{T(X)\oplus N^{\prime}}.$
In particular, just as in Remark 3.4, we find $G_{i}=A_{T(X)}$. Note, moreover, that $A_{T(X)}\subset A_{N^{\prime}}$ is the unique subgroup of $A_{N^{\prime}}$ order $\operatorname{disc}(T(X))$. Therefore, the isometry $\gamma_{i}\colon A_{T(X)}\simeq A_{N^{\prime}}$ has image $A_{T(X)}$ and therefore induces an isometry $A_{T(X)}\simeq A_{T(X)}$, which we shall also denote by $\gamma_{i}$.
We claim that the following is a well-defined bijection:
[…]
Let us begin by showing that $\Gamma$ is well-defined and injective. If $i,i^{\prime}\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ are two primitive embeddings contained in $\operatorname{Emb}_{N^{\prime}}(T(X),\Lambda_{\text{cub}}^{0})$. Then, by Lemma 2.4, $i$ and $i^{\prime}$ are in the same $O_{\mathrm{Hodge}}(T(X))$-orbit, i.e. there is a commutative square
${T(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$$\scriptstyle{\psi}\scriptstyle{\simeq}$${\Lambda_{\text{cub}}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{\phi}$${T(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i^{\prime}}$${\Lambda_{\text{cub}}^{0},}$
where $\psi\in O_{\mathrm{Hodge}}(T(X))$, if and only if there is an isometry $\psi^{\prime}\colon N^{\prime}\simeq N^{\prime}$ fitting into a commutative diagram
${G_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma_{i}}$$\scriptstyle{\overline{\psi}}$${A_{N^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{\psi^{\prime}}}$${G_{i^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma_{i^{\prime}}}$${A_{N^{\prime}}.}$
This exactly means that $\gamma_{i}$ and $\gamma_{i^{\prime}}$ define the same element of $O(N^{\prime})\backslash O(A_{T(X)})/O_{\mathrm{Hodge}}(T(X)),$ hence the map $\Gamma_{N^{\prime}}$ is well-defined and injective.
For surjectivity, note that any isometry $\gamma\colon A_{T(X)}\simeq A_{T(X)}$ induces the pair $(A_{T(X)},\gamma)$. From this pair, we obtain a primitive embedding $T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ with orthogonal complement isometric to $N^{\prime}$ as in Remark 2.3. This shows that $\Gamma_{N^{\prime}}$ is surjective.
The result now follows from Corollary 3.11. ∎
Due to the symmetry of the situation, one might expect an similar counting formula for Fourier–Mukai partners of cubic fourfolds to hold when we assume $\operatorname{disc}(A(X)_{\mathrm{prim}})$, rather than $\operatorname{disc}(T(X))$, is not divisible by $3$. The fundamental difference between these two assumptions is the following. If $X$ is a cubic fourfold such that $\operatorname{disc}(T(X))$ is not divisible by $3$, then $\operatorname{disc}(T(Y))$ is also not divisible by $3$ for any Fourier–Mukai partner $Y$ of $X$. On the other hand, if we assume $\operatorname{disc}(A(X)_{\mathrm{prim}})$ to be indivisible by $3$, then $X$ may have several Fourier–Mukai partners $Y$ for which $\operatorname{disc}(A(Y)_{\mathrm{prim}})$ is a multiple of $3$. Nevertheless, we have the following Counting Formula for Fourier–Mukai partners of $X$ whose primitive algebraic lattice has discriminant not divisible by $3$.
Proposition 3.12.
Let $X$ be a cubic fourfold which satisfies Derived Torelli, and write $N\coloneqq A(X)_{\mathrm{prim}}$. Suppose that $\operatorname{disc}(N)$ is not divisible by $3$. Then for any $N^{\prime}\in\mathcal{H}(N)$, we have
(3.3) $\operatorname{FM}(X,N^{\prime})=|O(N^{\prime})\backslash O(A_{N})/O_{\mathrm{Hodge}}(T(X))|.$
Here, the action of $O_{\mathrm{Hodge}}(T(X))$ on $O(A_{N})$ is the one defined by Remark 3.4.
Proof.
We fix an isometry $A_{T(X)}\simeq A_{N^{\prime}}\oplus C_{3}$, where $C_{3}$ is one of the two non-degenerate finite quadratic modules of order $3$. A primitive embedding $i\colon T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ induces a pair $(G_{i},\gamma_{i})$, where $G_{i}\subset A_{T(X)}$ and $\gamma_{i}\colon G_{i}\hookrightarrow A_{N^{\prime}}$. The group $G_{i}$ is equal to $A_{N^{\prime}}$ as a subgroup of $A_{T(X)}$, since $A_{N^{\prime}}\subset A_{T(X)}$ is the unique maximal subgroup whose order is not divisible by $3$. Therefore, $\gamma_{i}$ can be viewed as an element of $O(A_{N^{\prime}}).$
Similarly to the proof of Theorem 3.5, we obtain a map
[…]
which we claim to be a bijection. The proof the bijectivity of $\Gamma_{N^{\prime}}$ is completely analogous to the proof of Theorem 3.5, so we omit it here. ∎
Remark 3.13.
In the setting of Proposition 3.12, it is possible that $X$ has a Fourier–Mukai partner $Y$ with $N^{\prime}=A(Y)_{\mathrm{prim}}\notin\mathcal{H}(N)$. This is equivalent to the discriminant of $A(Y)_{\mathrm{prim}}$ being divisible by $3$. In this case, the gluing group of the embedding $N^{\prime}\hookrightarrow H^{4}(Y,\mathbb{Z})_{\mathrm{prim}}$ is isomorphic to $A_{T(Y)}$, but it is no longer necessarily the case that $A_{N^{\prime}}$ can be written as $A_{T(Y)}\oplus C_{3}$ for some finite quadratic form $C_{3}$ of order $3$. It is still possible to compute $\operatorname{FM}(X,N^{\prime})$ via Proposition 3.10, c.f. [9], but a simple formula such as (3.3) seems to be unattainable in this case. See Remark 4.8 for examples of this phenomenon.
4. Examples of $\mathbb{L}$-equivalent cubic fourfolds
In this section, we construct new examples of $\mathbb{L}$-equivalent cubic fourfolds, using the counting formula for Fourier–Mukai partners of Section 3. These examples are based on the following result.
Proposition 4.1.
[26, §5] [31, §2.6.1] Let $X$ be a cubic fourfold with one of the following properties:
  • A)
    $X\in\mathcal{C}_{14}$, i.e. $X$ is a Pfaffian cubic fourfold
  • B)
    $X$ contains a plane $P$ and another cycle $T\in\operatorname{CH}^{2}(X)$ such that the intersection number $(h^{2}-P)\cdot T$ is odd
Then there exists a K3 surface $S$ such that there exists a Hodge isometry $T(S)(-1)\simeq T(X)$ and such that the class of $X$ in the Grothendieck ring of varieties is given by
$X]=1+\mathbb{L}^{2}+\mathbb{L}^{4}+[S]\mathbb{L}.$
In light of Proposition 4.1, Pfaffian cubic fourfolds seem like a good place to start looking for $\mathbb{L}$-equivalent cubic fourfolds. However, very general Pfaffian cubic fourfolds have no Fourier–Mukai partners. This result follows immediately from [13, Proposition 2.6], and also from [51, Theorem 1.1] combined with [46, Proposition 1.10]. Here, we give a proof using Theorem 3.5.
Proposition 4.2.
Let $X$ be a cubic fourfold that is a very general member of $\mathcal{C}_{14}$, that is, such that $\operatorname{rk}A(X)=2$. Then $\operatorname{FM}(X)=\{X\}$.
Proof.
Since $A(X)\simeq K_{14}$ has rank $2$, it follows that $\operatorname{rk}A(X)_{\text{prim}}=1$. Therefore, $N\coloneqq A(X)_{\text{prim}}$ is clearly unique in its genus. Moreover, we have $\operatorname{disc}(T(X))=14$, and $A_{T(X)}$ is cyclic, since it is naturally a subgroup of the cyclic group $A_{N}$ by Remark 3.4. This implies that $A_{T(X)}$ has precisely $2$ isometries, namely $\pm\operatorname{id}_{A_{T(X)}}$. Indeed, if $x\in A_{T(X)}$ is a generator, then for any $\sigma\in O(A_{T(X)})$, we must have $\sigma(x)=ax$ for some $a\in(\mathbb{Z}/14\mathbb{Z})^{\times}$. However, since $x^{2}=\sigma(x)^{2}=a^{2}x^{2}\pmod{2\mathbb{Z}}$, it follows that we must have $a^{2}\equiv 1\pmod{14}$. Since $\pm 1$ are the only elements of $(\mathbb{Z}/14\mathbb{Z})^{\times}$ whose square equals $1$, we find $O(A_{T(X)})=\{\pm\operatorname{id}_{A_{T(X)}}\}$. Since both of these isometries are in the image of the group homomorphism $O_{\mathrm{Hodge}}(T(X))\to O(A_{T(X)})$, it follows from Theorem 3.5 that $\operatorname{FM}(X)=1$. ∎
Similarly, one might expect examples of $\mathbb{L}$-equivalent cubic fourfolds when studying cubic fourfolds containing two disjoint planes. If $X$ is a cubic fourfold containing two planes, which is very general with this property, then the algebraic part $A(X)$ of $H^{4}(X,\mathbb{Z})$ is isometric to the lattice
(4.1) $\left(\begin{matrix}3&1&1\\ 1&3&0\\ 1&0&3\end{matrix}\right)$
Remark 4.3.
Some of the lattice theoretical computations involving the next examples were done with OSCAR [49, 10]. We briefly recall here the behaviour of the key functions we used.
  • primitive_embeddings(L, N) (developed by Stevell Muller): computes primitive embeddings of the lattice N into L (optionally with classification data such as orthogonal complements).
  • vectors_of_square_and_divisibility(L, m, d) (developed by Stevell Muller): returns vectors $v\in L$ with $v^{2}=m$ and $\operatorname{div}(v)=d$.
  • genus_representatives(L) (developed by Simon Brandhorst and Stevell Muller): returns a set of lattices representing the genus of L.
The complete OSCAR code for checking the data of this paper can be found in the file ancillary to this paper.
Proposition 4.4.
Let $X$ be a cubic fourfold whose algebraic lattice $A(X)$ is isometric to (4.1). Then $\operatorname{FM}(X)=\left\{X\right\}$.
Proof.
Since $\operatorname{disc}(T(X))=\operatorname{disc}(A(X))=21$ is divisible by $3$, Theorem 3.5 is not applicable in this situation. Therefore, we use the counting algorithm of [9].
First, we determine that there is a unique lattice $N$ that arises as an orthogonal complement of an embedding $T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$. Explicitly, this lattice is given by
$N\coloneqq A(X)_{\text{prim}}\simeq\left(\begin{matrix}12&-3\\ -3&6\end{matrix}\right).$
Since $\operatorname{disc}(N)=63$, it follows from Lemma 2.5 that the gluing group $G=A_{T(X)}$, hence $A_{N}\simeq A_{T(X)}(-1)\oplus C_{3}$, where $C_{3}$ is one of the two finite quadratic modules of order $3$. Following the same reasoning as in the proof of Theorem 3.5, we obtain
$\operatorname{FM}(X)=|O(N)\backslash O(A_{T(X)})/O_{\mathrm{Hodge}}(T(X))|,$
where the action of $O(N)$ on $O(A_{T(X)})$ is the one defined in Remark 3.4. The natural map $O(N)\to O(A_{N})$ is surjective. Since $O(A_{N})\simeq O(A_{T(X)}(-1))\oplus O(C_{3})$, it follows that $O(N)$ acts transitively on $O(A_{T(X)})$. Therefore, we have $\operatorname{FM}(X)=1$, as required. ∎
We now turn our attention to the lattices of the form
$L_{n}\coloneqq\left(\begin{matrix}3&1&1\\ 1&3&0\\ 1&0&n\end{matrix}\right),$
where $n\geq 1$. If $X$ is a cubic fourfold with $A(X)\simeq L_{n}$, then $X$ satisfies property B of Proposition 4.1, so we are able to compute the class of $X$ in $K_{0}(\operatorname{Var}_{\mathbb{C}})$.
Importantly, if $(S_{1}$, $S_{2}$, $S_{3})$ is an ordered $\mathbb{Z}$-basis for $A(X)$ whose associated Gram matrix is $L_{n}$, then $S_{1}$ is the square of a hyperplane class if and only if $n$ is odd. On the other hand, if $n$ is even, then $S_{2}$ is the square of a hyperplane class.
By Proposition 4.4, $X$ has no Fourier–Mukai partners when $n=3$. Also, since the discriminant of $L_{n}$ is $\operatorname{disc}(L_{n})=8n-3$, we may use Theorem 3.5 to compute the number of Fourier–Mukai partners of $X$ whenever $n$ is not divisible by $3$.
Proposition 4.5.
Let $X$ be a cubic fourfold such that $A(X)\simeq L_{10}$. Assume moreover that $O_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}/2\mathbb{Z}$. Then $\operatorname{FM}(X)=2$. Moreover, the unique non-trivial Fourier–Mukai partner $Y$ of $X$ has $A(Y)_{\text{prim}}\simeq A(X)_{\text{prim}}$. In addition, $X$ and $Y$ are trivially $\mathbb{L}$-equivalent.
Proof.
Since $\operatorname{disc}(T(X))=\operatorname{disc}(L_{10})=77$ is not divisible by $3$, we may compute $\operatorname{FM}(X)$ using Theorem 3.5. A direct computation reveals
$N\coloneqq A(X)_{\text{prim}}\simeq\left(\begin{matrix}24&-3\\ -3&10\end{matrix}\right).$
This lattice is not unique in its genus. Indeed, we have $\mathcal{G}(N)=\left\{N,N^{\prime}\right\}$ where
$N^{\prime}\coloneqq\left(\begin{matrix}6&-3\\ -3&40\end{matrix}\right).$
Note that $N^{\prime}$ contains a vector $v$ with $v^{2}=6$ and $\operatorname{div}(v)=3$. Therefore, $\mathcal{H}(N)=\left\{N\right\}$. Since $A_{T(X)}$ has order $77$, it follows that $O(A_{T(X)})\simeq\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z}$. However, we have $O(N)\simeq\mathbb{Z}/2\mathbb{Z}$, hence
$|O(N)\backslash O(A_{T(X)})/O_{\mathrm{Hodge}}(T(X))|=2.$
That is, by Theorem 3.5, $X$ has a unique Fourier–Mukai partner $Y$, and we have $A(Y)_{\text{prim}}\simeq A(X)_{\text{prim}}\simeq N$.
It remains to be shown that $X$ and $Y$ are $\mathbb{L}$-equivalent. To prove this, we argue that $A(X)\simeq A(Y)$. Indeed, since $T(X)\simeq T(Y)$, and $A(X)=T(X)^{\perp}\subset H^{4}(X,\mathbb{Z})$, it follows that $A(X)$ and $A(Y)$ are in the same genus. However, the genus of $L_{10}$ consists of 5 lattices. Since $N\oplus\langle 3\rangle\simeq A(Y)_{\text{prim}}\oplus\langle 3\rangle\subset A(Y)$, contains both a vector of square $10$ and a vector of square $3$. Of the 5 lattices in $\mathcal{G}(L_{10})$, only $L_{10}$ and the lattice
$L_{10}^{\prime}\coloneqq\left(\begin{matrix}3&-1&0\\ -1&4&0\\ 0&0&7\end{matrix}\right)$
contain both a vector of square $10$ and a vector of square $3$. However, $L_{10}^{\prime}$ contains exactly 1 vector of square $3$, and its orthogonal complement is
$\left(\begin{matrix}7&0\\ 0&33\end{matrix}\right),$
which is not an even lattice. However, since $A(Y)_{\text{prim}}\subset A(Y)$ is an even lattice which is the orthogonal complement of a vector of square $3$, $A(Y)$ cannot be isometric to $L_{10}^{\prime}$ and thus it must be isometric to $L_{10}$. In particular, both $X$ and $Y$ satisfy property B of Proposition 4.1, hence there exist K3 surfaces $S,S^{\prime}$ such that
$X]=1+\mathbb{L}^{2}+\mathbb{L}^{4}+[S]\mathbb{L},\qquad[Y]=1+\mathbb{L}^{2}+\mathbb{L}^{4}+[S^{\prime}]\mathbb{L}.$
Moreover, we also obtain a Hodge isometry $T(S)(-1)\simeq T(X)\simeq T(Y)\simeq T(S^{\prime})(-1)$ from Proposition 4.1. In particular $S$ and $S^{\prime}$ are Fourier–Mukai partners. However, we compute that
$\operatorname{NS}(S)\simeq\left(\begin{matrix}2&7\\ 7&-14\end{matrix}\right),$
and one can compute using [19, Theorem 2.3] that $S$ has no non-trivial Fourier–Mukai partner, hence $S\simeq S^{\prime}$. In conclusion, we have
$X]=[Y]\in K_{0}(\operatorname{Var}_{\mathbb{C}}),$
so that $X$ and $Y$ are trivially $\mathbb{L}$-equivalent. ∎
Remark 4.6.
The positive integer $n=10$ is the smallest for which the lattice $L_{n}$ gives an example. The next ones are $n=11,17,26,28$. A complete classification is not known, but may be possible with the same techniques, as well as experimenting with other lattices. It would also be interesting to give a geometric description of such cubic fourfolds.
We now turn our attention to Pfaffian cubic fourfolds, and consider the lattice
$L_{a,b}\coloneqq\left(\begin{matrix}14&a\\ a&2b\end{matrix}\right),$
where $a,b\in\mathbb{Z}$ are chosen in such a way that $L_{a,b}$ is a non-degenerate positive-definite lattice. We will also write $A_{a,b}\coloneqq A_{L_{a,b}}.$
Proposition 4.7.
Let $X$ be a cubic fourfold such that $A(X)\simeq L_{2,9}$. Assume moreover that $O_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}/2\mathbb{Z}$. Then $\operatorname{FM}(X,L_{2,9})=2$. That is, $X$ has a unique Fourier–Mukai partner $Y$ with the property that $A(Y)_{\mathrm{prim}}\simeq A(X)_{\mathrm{prim}}$. Moreover, $X$ and $Y$ are trivially $\mathbb{L}$-equivalent.
Proof.
Since $\operatorname{disc}(L_{2,9})=248$ is not divisible by $3$, we may use Proposition 3.12 to compute
$\operatorname{FM}(X,L_{2,9})=|O(L_{2,9})\backslash O(A_{2,9})/O_{\mathrm{Hodge}}(T(X))|.$
We find $O(A_{2,9})\simeq\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z}$ and $O(L_{2,9})\simeq\mathbb{Z}/2\mathbb{Z}$. Thus, since $O_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}/2\mathbb{Z}$ by assumption, we obtain
$\operatorname{FM}(X,L_{2,9})=2.$
Denote by $Y$ the unique Fourier–Mukai partner of $X$ with $A(Y)_{\mathrm{prim}}\simeq L_{2,9}$. Then we have $Y\in\mathcal{C}_{14}$, i.e. $Y$ is a Pfaffian cubic fourfold and thus satisfies Property A of Proposition 4.1. In particular, we find K3 surfaces $S$, $S^{\prime}$ such that
$X]=1+\mathbb{L}^{2}+\mathbb{L}^{4}+[S]\mathbb{L},\qquad[Y]=1+\mathbb{L}^{2}+\mathbb{L}^{4}+[S^{\prime}]\mathbb{L}.$
Moreover, $S$ and $S^{\prime}$ are Fourier–Mukai partners. One can compute that
$\operatorname{NS}(S)\simeq\left(\begin{matrix}26&22\\ 22&-10\end{matrix}\right),$
but such a K3 surface never has non-trivial Fourier–Mukai partners, so we have $S\simeq S^{\prime}$, hence $X$ and $Y$ are trivially $\mathbb{L}$-equivalent. ∎
Remark 4.8.
A cubic fourfold $X$ as in Proposition 4.7 can be shown to have more Fourier–Mukai partners. For example, the genus of $L_{2,9}$ consists of three lattices:
$L_{2,9}=\left(\begin{matrix}14&2\\ 2&18\end{matrix}\right),\quad N_{1}\coloneqq\left(\begin{matrix}126&16\\ 16&4\end{matrix}\right)\simeq\left(\begin{matrix}62&0\\ 0&4\end{matrix}\right),\quad N_{2}\coloneqq\left(\begin{matrix}126&2\\ 2&2\end{matrix}\right).$
However, $N_{2}$ clearly contains a vector of square $2$, hence $N_{2}\notin\mathcal{H}(L_{2,9})$, and $X$ has no Fourier–Mukai partner $X^{\prime}$ with $A(X^{\prime})_{\mathrm{prim}}\simeq N_{2}$. On the other hand, $N_{1}$ contains neither vectors of square 2 nor vectors of square 6 and divisibility 3. One computes using Proposition 3.12 that $X$ has precisely $1$ Fourier–Mukai partner $X^{\prime}$ with $A(X^{\prime})_{\mathrm{prim}}\simeq N_{1}$.
Moreover, there is a primitive embedding $T(X)\hookrightarrow\Lambda_{\text{cub}}^{0}$ whose orthogonal complement is isometric to the lattice
$N_{3}\coloneqq\left(\begin{matrix}66&12\\ 12&36\end{matrix}\right).$
Importantly, $N_{3}$ has discriminant $2242$, which is divisible by $3$. Since $N_{3}$ contains no vectors of square $6$ nor vectors of square $2$, it follows that $X$ has at least one Fourier–Mukai partner $X^{\prime}$ with $A(X^{\prime})_{\mathrm{prim}}\simeq N_{3}$. Finally, the genus of $N_{3}$ consists of $N_{3}$ and the lattice
$N_{4}\coloneqq\left(\begin{matrix}18&-60\\ -60&324\end{matrix}\right),$
which again contains no vectors of square $6$ or $2$, hence $X$ has at least one Fourier–Mukai partner corresponding to $N_{4}$ as well.
It is an interesting question whether the Fourier–Mukai partners of $X$ whose primitive algebraic lattices are not isometric to that of $X$ are $\mathbb{L}$-equivalent to $X$.
Remark 4.9.
The techniques of the proof of Proposition 4.7 can be used to prove that the next lattices which generate examples of trivially $\mathbb{L}$-equivalent Pfaffian cubic fourfolds in precisely the same way as Proposition 4.7 are
$L_{4,6},\quad L_{5,6},\quad L_{6,5},\quad L_{6,10},\quad L_{8,6},\quad L_{9,8}.$
5. Implications of $\mathbb{L}$-equivalence
The following proposition is the analogue for cubic fourfolds of [11, Proposition 3.3]. We remark that for cubic fourfolds which satisfy Derived Torelli, this follows immediately from [21, Theorem 1.1].
Proposition 5.1.
Let $X$ be a cubic fourfold. Then there exist finitely many isomorphism classes of cubic fourfolds $Y$ such that there is an isomorphism (but not necessarily an isometry) of Hodge structures $f\colon H^{4}(X,\mathbb{Z})\to H^{4}(Y,\mathbb{Z})$.
Proof.
Passing to Hodge structures of weight $2$ we set $H_{X}:=H^{4}(X,\mathbb{Z})(1)$ and $H_{Y}:=H^{4}(Y,\mathbb{Z})(1)$. We have induced isomorphisms of algebraic and transcendental lattices (see Definition 2.8):
$A(X)=\operatorname{NS}(H_{X})\cong\operatorname{NS}(H_{Y})=A(Y),\qquad T(X)(1)=T(H_{X})\cong T(H_{Y})=T(Y)(1).$
In particular, by Lemma 2.18 we have also an equality $D:=\operatorname{disc}(T(X))=\operatorname{disc}(T(Y))$.
Denote by $S_{D}$ the set of all symmetric pairings $T(H_{X})\otimes T(H_{X})\to\mathbb{Z}(-2)$ which are morphisms in $\operatorname{HS}_{\mathbb{Z},4}$ and have discriminant $D$. The isomorphism $f(1)$ pulls back the intersection form on $H^{4}(Y,\mathbb{Z})$, and restricting to the transcendental part gives an element of $S_{D}$.
Consider the endomorphism ring $R:=\operatorname{End}_{\operatorname{HS}_{\mathbb{Z},2}}(T(H_{X}))$, and the group $G:=\operatorname{Aut}_{\operatorname{HS}_{\mathbb{Z},2}}(T(H_{X}))$ with its natural action on $S_{D}$. Up to Hodge automorphisms of $T(H_{X})$, the possible transcendental intersection forms coming from such $Y$ are parametrized by the $G$-orbits in $S_{D}$.
Now we need to show that the number of $G$-orbits in $S_{D}$ are finite. We follow the same proof as in [11, Proposition 3.3]. The only change is that we use $T(H_{X})$ instead of the transcendental lattice of a K3 surface. For carrying on the proof we need to check that $T(H_{X})$ satisfies the hypotheses of Zarhin’s theorem [55] (namely that $T(H_{X})$ is a polarized weight $2$ Hodge structure of K3 type). As a consequence, the $\mathbb{Q}$-algebra $R\otimes\mathbb{Q}$ is a totally real number field, or an imaginary quadratic extension of a totally real number field. Denote by $\beta$ an element of $S_{D}$ and consider the associated Hodge endomorphism $a(\beta)\in F$. Efimov’s argument then applies verbatim and gives the finiteness of $S_{D}/G$.
Finally, for each of these finitely many orbits we can apply Lemma 5.2 and get finitely many isomorphism classes of cubic fourfold $Y$, concluding the proof. ∎
Lemma 5.2.
Let $X$ be a cubic fourfold. Then there are exist finitely many isomorphism classes of cubic fourfolds $Y$ such that there is a Hodge isometry between $H^{4}(X,\mathbb{Z})$ and $H^{4}(Y,\mathbb{Z})$.
Proof.
Assume first we have a Hodge isometry $\phi\colon H^{4}(X,\mathbb{Z})\to H^{4}(Y,\mathbb{Z})$, and consider the class $v\coloneqq\phi^{-1}(h_{Y}^{2})$. Since $\phi$ is a Hodge morphism and $h_{Y}^{2}\in A(Y)$, we have $v\in A(X)$, and since $\phi$ is an isometry $(v,v)=(h_{Y}^{2},h_{Y}^{2})=3$. The lattice $A(X)$ is positive definite. In a positive definite integral lattice, there are only finitely many vectors of a given square, see [16, Section 2.1] for a reference. Hence the set $S_{3}:=\{v\in A(X):(v,v)=3\}$ is finite. Fix $v\in S_{3}$ and assume $Y_{1}$ and $Y_{2}$ are cubic fourfolds with Hodge isometries $\phi_{i}\colon H^{4}(X,\mathbb{Z})\to H^{4}(Y_{i},\mathbb{Z})$ such that $\phi_{i}^{-1}(h_{Y_{i}}^{2})=v$. The composition $\phi_{2}\circ\phi_{1}^{-1}$ is a Hodge isometry between $H^{4}(Y_{1},\mathbb{Z})$ and $H^{4}(Y_{2},\mathbb{Z})$. Moreover, this composition sends $h_{1}^{2}$ to $h_{2}^{2}$ and so gives an identification $H^{4}(Y_{1},\mathbb{Z})_{\operatorname{prim}}\cong H^{4}(Y_{2},\mathbb{Z})_{\operatorname{prim}}$. By Theorem 2.9 we have $Y_{1}\cong Y_{2}$.
Since $S_{3}$ is finite, there are only finitely many isomorphism classes of cubic fourfolds $Y$ admitting a Hodge isometry $H^{4}(X,\mathbb{Z})\cong H^{4}(Y,\mathbb{Z})$. ∎
The following theorem is the analogue for cubic fourfolds of [11, Theorem 3.4]. Again, for cubic fourfolds which satisfy Derived Torelli, Theorem 5.4 implies that $\mathbb{L}$-equivalent cubic fourfolds are Fourier-Mukai partners, and the results again follows from [21, Theorem 1.1].
Theorem 5.3.
Let $X$ be a cubic fourfold. Then there exist finitely many cubic fourfolds $Y$ that are $\mathbb{L}$-equivalent to $X$.
Proof.
Let $Y$ be a cubic fourfold such that $X\sim_{\mathbb{L}}Y$, and consider the integral Hodge realisation $\operatorname{Hdg}_{\mathbb{Z}}$ constructed in [11, Section 3]. From $[X]=[Y]$ in $K_{0}(\operatorname{Var}_{\mathbb{C}})[\mathbb{L}^{-1}]$ we deduce $H^{4}(X,\mathbb{Z})]=[H^{4}(Y,\mathbb{Z})]$ in $K_{0}(\operatorname{HS}^{4}_{\mathbb{Z}})$.
By [11, Theorem 2.7], applied to the additive category $\operatorname{HS}^{4}_{\mathbb{Z}}$, the class $H^{4}(X,\mathbb{Z})]$ in $K_{0}(\operatorname{HS}^{4}_{\mathbb{Z}})$ contains only finitely many isomorphism classes of objects. In particular there are only finitely many possibilities for $H^{4}(Y,\mathbb{Z})$ as an integral Hodge structure when $Y$ ranges over cubic fourfolds with $X\sim_{\mathbb{L}}Y$. The thesis then follows from Proposition 5.1. ∎
The following result appears as Theorem 1.4 in [4].
Theorem 5.4.
Let $X$ and $Y$ be cubic fourfolds which satisfy Derived Torelli (for example, if $X$ is a cubic fourfold satisfying one of the properties of Proposition 3.2) Moreover, assume $\operatorname{rk}(T(X))\neq 4$ and $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, if $X$ and $Y$ are $\mathbb{L}$-equivalent, they are Fourier–Mukai partners.
Proof.
By Theorem 2.20, there is a Hodge isometry $T(X)\cong T(Y)$. Then, if $X$ and $Y$ satisfy Derived Torelli, then they are Fourier–Mukai partners. ∎
The following result appears as Theorem 1.2 in [4].
Proposition 5.5.
Let $X$ be a cubic fourfold not contained in any of the Hassett divisors. Assume moreover that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then, for any cubic fourfold $Y$, we have
$X\sim_{\mathbb{L}}Y\iff X\simeq Y$
Proof.
Note that $\operatorname{rk}(T(X))=22\neq 4$. Therefore, we may apply Theorem 2.20 to obtain a Hodge isometry $T(X)\simeq T(Y)$. However, since $X$ and $Y$ are not contained in any Hassett divisor, we have
$H^{4}(X,\mathbb{Z})_{\operatorname{prim}}=T(X)\simeq T(Y)=H^{4}(Y,\mathbb{Z})_{\operatorname{prim}},$
hence $X\simeq Y$ by the Torelli Theorem 2.9. ∎
The following result appears as Proposition 4.4 in [4].
Proposition 5.6.
Let $X$ and $Y$ be cubic fourfolds. Suppose $X$ and $Y$ are $\mathbb{L}$-equivalent. Then their Fano varieties of lines are also $\mathbb{L}$-equivalent.
Proof.
Since $X$ and $Y$ are assumed to be $\mathbb{L}$-equivalent, $X^{[2]}$ and $Y^{[2]}$ are $\mathbb{L}$-equivalent by [47, Lemma 2.1]. Now the result follows directly from the beautiful formula of Galkin–Shinder [14]: for any cubic fourfold $X$, we have
$\mathbb{L}^{2}[F(X)]=[X^{[2]}]-[\mathbb{P}^{4}][X].$
Theorem 5.7.
Let $X$ and $Y$ be cubic fourfolds. Assume that $X$ satisfies Derived Torelli, that $\operatorname{rk}(T(X))\neq 4$, and that $\operatorname{End}_{\mathrm{Hodge}}(T(X))\simeq\mathbb{Z}$. Then if $F(X)\sim_{\mathbb{L}}F(Y)$, there is an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$, hence also an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$.
Proof.
By [41, Theorem 3.5] and Lemma 2.19, the gluing group
$G=\frac{H^{2}(F(X),\mathbb{Z})}{T(F(X))\oplus\operatorname{NS}(F(X))}$
has order $|G|=\operatorname{disc}(T(F(X)))$. Therefore, if $F(X)$ is $\mathbb{L}$-equivalent to $F(Y)$, it follows from Lemma 2.17 that $\operatorname{disc}(T(F(X)))=\operatorname{disc}(T(F(Y)))$. Since $T(F(X))$ is Hodge anti-isometric to $T(X)$, we find that $[T(X)(1)]=[T(Y)(1)]$. By Theorem 2.20, there is a Hodge isometry $T(X)\simeq T(Y)$, hence an equivalence $\mathcal{A}_{X}\simeq\mathcal{A}_{Y}$ by Definition 1.7. The fact that there is an equivalence $\mathcal{D}^{b}(F(X))\simeq\mathcal{D}^{b}(F(Y))$ follows from the main result of [30]. ∎
The proof of the following result was communicated to us by Evgeny Shinder.
Proposition 5.8.
Let $X$ and $Y$ be projective hyperkähler manifolds. Suppose $X$ and $Y$ are birational. Then we have $X]=[Y]\in K_{0}(\operatorname{Var}_{\mathbb{C}})$.
Proof.
By [22, Theorem 4.6] there exist two smooth projective families $\mathcal{X}$ and $\mathcal{Y}$ over a smooth one dimensional base $S$ with special fibres $\mathcal{X}_{0}\cong X$ and $\mathcal{Y}_{0}\cong Y$, such that $\mathcal{X}$ and $\mathcal{Y}$ are isomorphic over $S\smallsetminus\{0\}$.
Now consider the generic fibres over $S\smallsetminus\{0\}$. As in [44] let $R$ be the completion of the local ring of $S$ at 0 and let $K=\operatorname{Frac}(R)$, so that $\operatorname{Spec}K$ is the generic point of $\operatorname{Spec}(R)$. By taking base changes we get the restrictions $\mathcal{X}_{R}$ and $\mathcal{Y}_{R}$ of $\mathcal{X}$ and $\mathcal{Y}$ to a formal neighbourhood of 0, and the corresponding generic fibres $\mathcal{X}_{K}$ and $\mathcal{Y}_{K}$.
Recall that we have a canonical identification of special fibres $(\mathcal{X}_{R})_{k}=\mathcal{X}_{R}\times_{\operatorname{Spec}R}\operatorname{Spec}k\cong\mathcal{X}\times_{S}\operatorname{Spec}k\cong\mathcal{X}_{0}$, and similarly $(\mathcal{Y}_{R})_{k}\cong\mathcal{Y}_{0}$.
The isomorphism of the families over $S\smallsetminus\{0\}$ induces an isomorphism over $K$ of $\mathcal{X}_{K}$ and $\mathcal{Y}_{K}$, which induces $\mathcal{X}_{K}]=[\mathcal{Y}_{K}]$ in $K_{0}(\operatorname{Var}_{K})$.
Now apply the motivic volume morphism $\operatorname{Vol}_{K}:K_{0}(\operatorname{Var}_{K})\to K_{0}^{\hat{\mu}}(\operatorname{Var}_{\mathbb{C}})$ of [44, Theorem 3.1.1]. From $\mathcal{X}_{K}]=[\mathcal{Y}_{K}]$ we get $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=\operatorname{Vol}_{K}([\mathcal{Y}_{K}])$. From the fact that $\mathcal{X}_{R}$ ($\mathcal{Y}_{R}$) is a smooth proper $R$ model of $\mathcal{X}_{K}$ ($\mathcal{Y}_{K}$), so $\operatorname{Vol}_{K}([\mathcal{X}_{K}])=[\mathcal{X}_{0}]=[X]$ and $\operatorname{Vol}_{K}([\mathcal{Y}_{K}])=[\mathcal{Y}_{0}]=[Y]$. This concludes the proof. ∎
We obtain the following as an immediate consequence:
Corollary 5.9.
Let $X$ and $Y$ be cubic fourfolds. If $F(X)$ is birational to $F(Y)$, then $F(X)$ and $F(Y)$ are $L$-equivalent. In other words, we have $\mathbf{BF}\implies\mathbf{LF}$.